OpenStudy (abbycross167):

Will someone please help me figure out how to answer this question? At Generic High School, in the year 2001, there were 20 students per every computer. In 2008, there were 4.5 students per computer. During which year were there 10 students per computer?

OpenStudy (anonymous):

is this relation supposed to be linear?

OpenStudy (anonymous):

if so, in 8 years it went from 20 to 4.5 making your slope \(\frac{4.5-20}{8}\)

OpenStudy (abbycross167):

Umm I really don't know... I think so

OpenStudy (abbycross167):

@satellite73

OpenStudy (anonymous):

\[\frac{4.5-20}{7}=-\frac{31}{14}\]

OpenStudy (anonymous):

your linear model is \[y=-\frac{31}{14}+20\] set that equal to \(10\) and solve for \(x\)

OpenStudy (anonymous):

that is wrong sorry it is\[y=-\frac{31}{14}x+20\]

OpenStudy (abbycross167):

17 11/14?

OpenStudy (abbycross167):

@jim_thompson5910 could you please try to help me answer the question?

jimthompson5910 (jim_thompson5910):

ok one moment

OpenStudy (abbycross167):

ok thank you

jimthompson5910 (jim_thompson5910):

use the equation @satellite73 wrote out plug in y = 10 and solve for x \[\Large y = -\frac{31}{14}x + 20\] \[\Large 10 = -\frac{31}{14}x + 20\] \[\Large 10-20 = -\frac{31}{14}x + 20-20\] \[\Large -10 = -\frac{31}{14}x\] do you see how to finish?

OpenStudy (abbycross167):

umm maybe... would I divide -31/14 by -10?

jimthompson5910 (jim_thompson5910):

or multiply both sides by the reciprocal of the fraction

OpenStudy (abbycross167):

ok would you mind showing me so I know i got it right?

jimthompson5910 (jim_thompson5910):

ok so we have this equation 

jimthompson5910 (jim_thompson5910):

multiply both sides by -14/31 this is the reciprocal of -31/14